OpenStudy (anonymous):

determine the values of x given for which f'(x)>0, given that f(x)= 1/(4+2x^2)^4

OpenStudy (amistre64):

that would be easier to do if you simply take and negate the exponent and work the reciprocal

OpenStudy (amistre64):

f(x) = (4 +2x^2)^(-4) f'(x) = -4(4 +2x^2)^(-5) 4x

OpenStudy (amistre64):

so when x = 0; is all I see for a zero

OpenStudy (amistre64):

-16x ----------- = 0 when x = 0 (4 +2x^2)^5

OpenStudy (anonymous):

Real solutions anyway.

OpenStudy (amistre64):

yes :) assumably

OpenStudy (amistre64):

when x = 1 we get a - value; so stick with values of x<0

OpenStudy (anonymous):

thanks!